Explain how the reactions of Mg and MgO with HCL could be used with the Heat of fusion of H2O to determine the heat of combustion of Mg
1
Expert's answer
2020-12-23T09:36:46-0500

To determine the heat of combustion of Mg you will measure the heats of two reactions:

a) Mg + 2 HCl → MgCl2 + H2(g)     ΔHa = measured 

b) MgO + 2 HCl → MgCl2 + H2O(l)     ΔHb = measured 

You also need a ΔH value for reaction c, which is from the literature.

c) H2(g) + 1/2 O2(g) → H2O(l)     ΔHc = −285.8 kJ 

Using these ΔH values and Hess's Law, you can determine the ΔH for magnesium combustion, where magnesium metal is burned in oxygen.

Mg + 1/2 O2 → MgO     ΔHd = ??? 

1. Record the mass of the clean, dry calorimeter to 0.01 g.

2. Put 50 mL of distilled water into the calorimeter.

3. Add 25 mL of approximately 2 M HCl.

4. Record the mass of this water and acid combination to 0.01 g.

5. Take about 0.10–0.15 g of magnesium ribbon and record its mass to 0.01 g.

6. Take the initial temperature of the acidic water in the calorimeter to 0.1°C.

7. Add the metal, stir constantly, and record the maximum temperature reached to 0.1°C.

8. Repeat this procedure substituting about 0.30g of magnesium oxide for the magnesium ribbon.

The goal is to determine the heat of reaction for the combustion of magnesium to form the oxide. Target: Mg + 1/2 O2 → MgO     ΔHrxn = ?

∆H1 + ∆Hcalorimeter + ∆Hchemical reaction =

m1cwater(Tfinal – T1) + ccalorimeter(Tfinal – T1) + ∆Hreaction = 0

Using ΔHa, ΔHb, and ΔHc , employ Hess' Law to calculate the heat of reaction for the magnesium combustion reaction.